Lane invested $14000 part at 15% and part at 2%. If the total interest at the end of the year is $1060 how much did she invest at 15%. My dear friends, if anyone can help I would be so grateful.
Answer by checkley71(8403)

LET X BE THE AMOUNT INVESTED @ 15% & 14000-X BE THE AMOUNT INVESTED @ 2%.
.15X+.02(14000-X)=1060
.15X+280-.02X=1060
.13X=1060-280
.13X=780
X=780/.13
X=6000 ANSWE RFOR AMOUNT INVESTED @ 15%.
PROOF
.15*6000+.02*8000=1060
900+160=1060
1060=1060